OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4> 20.3° 10.2° 0.2° 30.3°

OpenStudy (anonymous):

I looked back into my lesson, I simply don't get how to convert this into degrees.

OpenStudy (astrophysics):

\[\cos \theta = \frac{ u \cdot v }{ ||u|| \times ||v|| }\]

OpenStudy (astrophysics):

Remember to take the arrcos though

OpenStudy (astrophysics):

The x is there just to multiply (to make it clear) it does NOT mean cross product

OpenStudy (anonymous):

I got it thanks!

OpenStudy (astrophysics):

Np :)